OpenStudy (anonymous):

Let y=5x2. Find the change in y, change in y when x=5 and change in x=0.4 Find the differential dy when x=5 and dx=0.4

OpenStudy (anonymous):

First of all, we'll have to subsitute x in the following equation to figure out the change in y and you'll get as follows:\[y = 5(5)^2 = 125\] so y = 125 when x = 5 as for the second one: \[y = 5(0.4)^2 = \] so y = 4/5 when x = 0.4 and the last part, differentiate the equation and you'll get : \[y = 10x\] I'll leave this one for you since I have showed you how to find the change of y in the first part, all you have to dp is substitute x to find out. Good Luck! ^_^

OpenStudy (anonymous):

do*